# Counting distinct states in physical dynamics

### v1

@@ Abstract
Maximum counts of distinct states define important physical quantities, the most famous being entropy. Here we show that energy, momentum and other basic quantities of mechanics are also maximum counts. This generalizes a property familiar in communication with classical waves: frequency-width equals maximum number of distinct values that can be sent, per unit time. Basic mechanical quantities are frequency-widths of wavefunction evolution, and similarly bound number of distinct states. Evolutions that achieve maximum distinctness are effectively discrete. These properties refine our understanding of uncertainty, measurement, classical mechanics, and relativity.
@@ Introduction
More general bounds on certainty, introduced here, directly count the maximum number of states in a quantum evolution that can be perfectly distinguished from each other. These distinctness bounds reveal, for example, that macroscopic energy and momentum are counts of distinct states—in time and space respectively. The prototype for the new bounds is Nyquist’s classical bound on communication with waves [26]. He showed that the critical resource that limits the number of distinct signal values that can be transmitted per unit time is bandwidth, the width of the range of frequencies that can appear in the wave’s Fourier decomposition. Intuitively, doubling bandwidth lets us double all frequencies, making everything happen twice as fast—including distinct values. In quantum evolutions, more general measures of frequency width similarly limit how fast things can change.
@@
The $E-E_{0}$ bound is particularly useful, since it identifies average energy as a conserved resource governing distinguishability [14] and hence measurability (cf. [31, 32, 33, 34, 35, 37, 36]). It also shows a system’s ground state energy cannot cause distinct state change within the system (cf. [38, 39, 40]).
@@
For any natural definition of energy width $h{\Delta\nu}$, we get a similar bound ${\Delta\nu}\,\tau\geq f$, where $f$ increases with $\Delta N$} but is always of order unity. The ${\Delta N}=2$ cases have a single interval $\tau$ and the smallest $f$. These special cases include known uncertainty bounds that treat time as a classical parameter (quantum speed limits) [14, 15, 13, 16, 12, 17, 18, 19].
@@
In all cases, ${\Delta\nu}\,\tau$ is minimized by a uniform energy distribution with $N$ frequencies $1/N\tau$ apart. This defines $f(N)$, which is of order one if $\Delta\nu$} is a natural width: a width that, for a uniform distribution, has the same order of magnitude as bandwidth $\nu_{\max}-\nu_{\min}$. As examples, we evaluate $f(N)$ for all generalized $M$th deviations from the mean or minimum (the widths are natural for $M\gtrsim 2/3$), and for a range of energies with total probability weight $q$ (the width is natural for $q>1/2$). The limit $N\to\infty$ of $f(N)$ bounds any evolution with a long distinct portion.
@@
Our other example, the width of an energy range $h\,\Delta\nu_{q}$ that holds probability $q$, is not a normal width. Periodic bound (17) can decrease with $\Delta N$} and so cannot apply exactly to portions. Figure 11 compares the best linear bound implied by (17), $\Delta\nu_{q}\tau\geq q-1/{\Delta N}$, to exact portion bounds (dots) determined numerically. The curved line is an achievable potion-bound for ${\Delta N}=2$ [17]:
@@
Other distinct evolutions
@@
Equality is obvious here for $t=k{\mkern 1.0mu}\tau$ with integer $k$, since $\operatorname{sinc}_{b}(k-n)=\delta_{kn}$. Since the sum is a superposition of the same Fourier components as $s(t)$, the discrete values at $t=k{\mkern 1.0mu}\tau$ determine all intermediate values, as we discussed in the introduction: $s(t)$ has both discrete and continuous properties [57, 58]. Center frequency $b/\tau$ affects phase of $\operatorname{sinc}_{b}(t/\tau{\mkern 1.0mu}-{\mkern 1.0mu}n)$ at intermediate $t$, but not at $t=k{\mkern 1.0mu}\tau$.
@@
Here we interpolate the evolution of each Hilbert-space component of the state vector $\left|{\Psi}(t)\right\rangle$ from its discrete values at the minimum time separation $\tau$ allowed by the bandwidth. This does not require the $\left|{\Psi}(n\tau)\right\rangle$ to be distinct, but if they are, $\left|{\Psi}(t)\right\rangle$ is a fastest changing (minimum bandwidth) distinct evolution. Then all frequencies used are equally weighted and the center frequency is average energy (over $h$). An infinite number of $N\gg 1$ distinctness bounds that give separation $\tau$ are all achieved by a single superposition (29) of distinct states $\left|{\Psi}(n\tau)\right\rangle$ (cf. [59, 60]). This analysis also applies to similar evolutions, such as shifts in space and rotations.
@@
and $\tau=T/N$. The limit ${\operatorname{sinc}_{b,\infty}}={\operatorname{sinc}_{b}}$. To have periodicity in both magnitude and phase, all frequencies must have a whole number of oscillations per period: for example, $\nu_{0}=0$ in (8). Thus $m$ must be an integer in (30), making this a periodic delta function for integer $u$: $\operatorname{sinc}_{b,\scriptscriptstyle N}u=1$ if $u\equiv 0\bmod N$, and 0 otherwise. Then, as before, we see from (28) that the $N$ discrete values at $t=k{\mkern 1.0mu}\tau$ suffice to determine the $N$ Fourier coefficients and hence $s(t)$, making it effectively discrete [61, 62]. Similarly from (29), $N$ states $\tau$ apart determine $\left|{\Psi}(t)\right\rangle$; if these states are distinct, evolution is maximally distinct.
With a half-integer number of oscillations per “period” (as in some rotations), the $\left|{\Psi}(n\tau)\right\rangle$ for $0{\mkern-1.0mu}\leq{\mkern-1.0mu}n{\mkern-1.0mu}\leq{\mkern-1.0mu}N{\mkern-1.0mu}-{\mkern-1.0mu}1$ still determine all Fourier coefficients in (29), but we gain a factor of 1-1 for each successive interval $N\tau$. For example, $b=0$, $N=2$ gives $\mbox{$\left|{\Psi}(t)\right\rangle$}=\cos{\frac{\pi t}{2\tau}}\,\mbox{$\left|{\Psi}(0)\right\rangle$}\,+\,\sin{\frac{\pi t}{2\tau}}\,\mbox{$\left|{\Psi}(\tau)\right\rangle$}$.
@@
Finite distinctness limits resolution in measurements. For example, suppose we want to measure the time between two states of a unitary evolution with average energy $E_{0}+\Delta E$. For maximum resolution the evolution should have as many distinct states as possible, to give a smallest possible spacing $\tau\approx h/\Delta E$ between distinguishable moments. Distinct states then form a discrete basis for the time evolution, as in (29), so the time between any two states is only defined up to a resolution $\tau$. In general, a minimum spacing between distinct “shifts” of a state limits the resolution of a measured shift (cf. [18]).
Since light is often used to probe systems of interest, the maximum density of distinct states in a periodic optical evolution is a key constraint on measurements. This is given by the energy bounds. Historically, resolution limits stem from a photon number bound on phase resolution in interferometry using monochromatic light [31, 32, 33, 34, 35, 37, 36]. If the light has frequency $1/T$, its time evolution is periodic with period $T$, and the average excitation energy is just the average number of photons $\left\langle n\right\rangle$ times $h/T$. Then (2) limits the number of distinct times to $N\leq 2\mbox{$\left\langle n\right\rangle$}+1$, so distinct phases are at least $2\pi/(2\mbox{$\left\langle n\right\rangle$}+1)$ apart. In general, though, photon number is not a surrogate for energy in bounding resolution in phase, since we can have the same excitation energy with one photon of frequency $k/T$ as with $k$ photons of frequency $1/T$. This lets us make a uniform superposition of single photon modes with $N$ different frequencies k/Tk/T—a single photon state—that traverses $N$ distinct states in period $T$ (see Figure 2).
@@
Maximum distinctness defines how distinct states are counted in classical phase space [41, 42]. Consider the number of distinct shifts allowed by the momentum bandwidth bound (25) for a large gas of conserved distinguishable particles: for each spatial degree of freedom, maximum distinctness allows one distinct state per area $h$ of accessible momentum-range $\times$ spatial-length. For identical particles, we divide by a permutation count.
@@
Figure 15: Most-distinct vs Gaussian. Fastest changing states under shifts (Figure 12) give a probability distribution $\mbox{$\left|\psi(x)\right|$}^{2}$ (solid) similar to same-height Gaussian distribution (dashed). Nevertheless, the standard deviation of $\mbox{$\left|\psi(x)\right|$}^{2}$ is infinite.
@@
If we model classical mechanics as maximally distinct, though, all evolution is effectively discrete. For large objects, discrete separations $h/2E$ in time can be arbitrarily small, but this is just a property of aggregated change: if enough clocks tick, average time between ticks can be arbitrarily short. Given the energy density $\rho$ in a region, we can identify the scale of individual “ticks” there. If we assume that, at the most microscopic scale, all motion is at the speed of light and all change is motion, we can estimate the distance $\lambda$ between distinct positions as a function of $\rho$. In natural units with $c=1$ and $h=2$, $\rho\,\lambda^{3}=E=p=1/\lambda$, since minimum distinct volume $\lambda^{3}$ must move a width $\lambda$ to move distinctly, so $\lambda=\rho^{-1/4}$. In the extreme case of $\rho$ the Planck density, $\lambda$ is the Planck length and $\tau=1/E=\lambda$ is the Planck time. Effective discreteness provides a fundamental link between classical spacetime and quantum dynamics (cf. [104, 105, 106]).
@@ Conclusions
Quantum mechanics generalizes Fourier analysis to describe all dynamics as unitary evolution of linear superpositions. Widths of generators of such evolutions play the role of bandwidth in bounding the number of distinct states possible in the course of the evolution. Taking $h$ as the unit of action, any natural width of the eigenvalue distribution, times the length of evolution, normally equals the maximum number of distinct states—up to a factor of order unity. For example, if $\Delta E$ is any deviation-width of the energy distribution for moment one or greater, at most about $\Delta E{\mkern 1.0mu}{\Delta t}$ distinct states can occur in time $\Delta t$}. Action similarly counts distinct states possible in a shift in space $(\Delta p_{x}{\mkern 1.0mu}\Delta x)$ or a rotation $(\Delta J_{\phi}\Delta\phi)$, since momentum is the generator of shifts and angular momentum the generator of rotations. Maximum distinctness implies discreteness: continuous evolution becomes just interpolation between distinct states. The number of distinct states possible in an evolution defines the resolution with which the length of the evolution can be measured.
@@
Acknowledgments   I thank Gerald J. Sussman and the MIT CSAIL Laboratory for aiding and encouraging this research. I also thank Sussman for valuable discussions, as well as Tom Toffoli, Ted Jacobson, Samuel Braunstein, Lorenzo Maccone and Adolfo del Campo.

### v2

@@ Abstract
A finite physical system has a finite entropy—a finite number of distinct possible states. Here we show that other basic quantities of physics are also finite counts of distinct states, including energy, momentum and Lagrangian action. A finite number of distinct states also limits the resolution of measurements and makes classical spacetime effectively discrete. Our analysis generalizes Nyquist’s classical bound on distinct signal values in waves that superpose a finite range of frequencies: distinct states of unitary transformations are bounded by more general widths of quantum superpositions. Finding the shortest transformations with $N$ distinct states also generalizes quantum speed limits.
@@ Introduction
More general bounds on certainty, introduced here, directly count the maximum number of states in a quantum evolution that can be perfectly distinguished from each other. These distinctness bounds reveal, for example, that macroscopic energy and momentum are counts of distinct states—in time and space respectively—and explain stationary action as a counting principle. The prototype for the new bounds is Nyquist’s classical bound on communication with waves [26]. He showed that the critical resource that limits the number of distinct signal values that can be transmitted per unit time is bandwidth, the width of the range of frequencies that can appear in the wave’s Fourier decomposition. Intuitively, doubling bandwidth lets us double all frequencies, making everything happen twice as fast—including distinct values. In quantum evolutions, more general measures of frequency width similarly limit how fast things can change.
@@
The $E-E_{0}$ bound is particularly useful, since it identifies average energy as a conserved resource governing distinguishability [14] and hence measurability (cf. [31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42]). It also shows a system’s ground state energy cannot cause distinct state change within the system (cf. [43, 44, 45]).
@@
For any natural definition of energy width $h{\Delta\nu}$, we get a similar bound ${\Delta\nu}\,\tau\geq f$, where $f$ increases with $\Delta N$} but is always of order unity. The ${\Delta N}=2$ cases have a single interval $\tau$ and the smallest $f$. These special cases include known uncertainty bounds that treat time as a classical parameter: quantum speed limits [14, 15, 13, 16, 12, 17, 18, 19].
@@
In all cases, ${\Delta\nu}\,\tau$ is minimized by a uniform energy distribution with $N$ frequencies $1/N\tau$ apart, and minima for all widths centered on the same frequency are achieved simultaneously. The uniform distribution defines $f(N)$, which is of order one if $\Delta\nu$} is a natural width: a width that, for a uniform distribution, has the same order of magnitude as bandwidth $\nu_{\max}-\nu_{\min}$. As examples, we evaluate $f(N)$ for all generalized $M$th deviations from the mean or minimum (the widths are natural for $M\gtrsim 2/3$), and for a range of energies with total probability weight $q$ (the width is natural for $q>1/2$). The limit $N\to\infty$ of $f(N)$ bounds any evolution with a long distinct portion.
@@
Our other example, the width of an energy range $h\,\Delta\nu_{q}$ that holds probability $q$, is not a normal width. Periodic bound (17) can decrease with $\Delta N$} and so cannot apply exactly to portions. Figure 11 compares the best linear bound implied by (17), $\Delta\nu_{q}\tau\geq q-1/{\Delta N}$, to exact portion bounds (dots) determined numerically. The curved line is an achievable portion-bound for ${\Delta N}=2$ [17]:
@@
Other distinct transformations
@@
Equality is obvious here for $t=k{\mkern 1.0mu}\tau$ with integer $k$, since $\operatorname{sinc}_{b}(k-n)=\delta_{kn}$. Since the sum is a superposition of the same Fourier components as $s(t)$, the discrete values at $t=k{\mkern 1.0mu}\tau$ determine all intermediate values, as discussed in the introduction. This also makes discrete integration of $s(t)$ exact [62, 63]. Center frequency $b/\tau$ affects phase of $\operatorname{sinc}_{b}(t/\tau{\mkern 1.0mu}-{\mkern 1.0mu}n)$ at intermediate $t$, but not at $t=k{\mkern 1.0mu}\tau$.
@@
Here we interpolate the evolution of each Hilbert-space component of state vector $\left|{\Psi}(t)\right\rangle$ from its discrete values at the minimum time separation $\tau$ allowed by bandwidth. This makes discrete integration of expectation values exact [64]. The identity (29) does not require the $\left|{\Psi}(n\tau)\right\rangle$ to be distinct, but if they are, $\left|{\Psi}(t)\right\rangle$ is a fastest changing (minimum bandwidth) distinct evolution. Then all frequencies used are equally weighted and center frequency $b/\tau$ is the average energy (over $h$). A superposition (29) of distinct $\left|{\Psi}(n\tau)\right\rangle$ achieves all distinctness bounds for $N\to\infty$ with the given $b$ and $\tau$ (cf. [65, 66]). This analysis also applies to other single parameter unitary transformations such as shifts in space and rotations.
@@
and $\tau=T/N$. The limit ${\operatorname{sinc}_{b,\infty}}={\operatorname{sinc}_{b}}$. To have periodicity in both magnitude and phase, all frequencies must have a whole number of oscillations per period: for example, $\nu_{0}=0$ in (8). Then $m$ must be an integer in (30), making this a periodic delta function for integer $u$: $\operatorname{sinc}_{b,\scriptscriptstyle N}u=1$ if $u\equiv 0\bmod N$, and 0 otherwise. Then, as before, we see from (28) that the $N$ discrete values at $t=k{\mkern 1.0mu}\tau$ suffice to determine the $N$ Fourier coefficients and hence $s(t)$, making it effectively discrete [67, 68]. Similarly from (29), $N$ states $\tau$ apart determine $\left|{\Psi}(t)\right\rangle$; if these states are distinct, evolution is maximally distinct.
With a half-integer number $m$ oscillations per “period” (as in some rotations), the $\left|{\Psi}(n\tau)\right\rangle$ for $0{\mkern-1.0mu}\leq{\mkern-1.0mu}n{\mkern-1.0mu}\leq{\mkern-1.0mu}N{\mkern-1.0mu}-{\mkern-1.0mu}1$ still determine all Fourier coefficients in (29), but we gain a factor of 1-1 for each successive interval $N\tau$. For example, $b=0$, $N=2$ gives $\mbox{$\left|{\Psi}(t)\right\rangle$}=\cos{\frac{\pi t}{2\tau}}\,\mbox{$\left|{\Psi}(0)\right\rangle$}\,+\,\sin{\frac{\pi t}{2\tau}}\,\mbox{$\left|{\Psi}(\tau)\right\rangle$}$.
@@
Finite distinctness limits resolution in measurements. For example, suppose we want to measure the time between two states of a unitary evolution with average energy $E_{0}+\Delta E$. For maximum resolution the evolution should have as many distinct states as possible to give the minimum spacing $\tau\approx h/2\Delta E$ between distinguishable moments. Distinct states then form a discrete basis for the time evolution, as in (29), so the time between any two states is only defined up to a resolution $\tau$. In general, a minimum spacing between distinct “shifts” of a state limits the resolution of a measured shift (cf. [18]).
Since light is often used to probe systems of interest, the maximum density of distinct states in a periodic optical evolution is a key constraint on measurements. This is given by the energy bounds. Historically, resolution limits stem from a photon number bound on phase resolution in interferometry using monochromatic light [31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42]. With light of frequency $\nu$, average energy bound $\tau\,\Delta E\geq h/4$ implies an average photon number $\bar{n}$ bound, since $\Delta E=\bar{n}h\nu$ and the period of time evolution $T\leq 1/\nu$. Thus $\bar{n}\,\theta\equiv\bar{n}\,(2\pi\,\tau{\mkern-1.0mu}{\mkern-1.0mu}/T)\geq 2\pi\tau\bar{n}\nu\geq\pi/2$. In general, though, only energy bounds distinctness in time.
For an ideal interferometer, maximally distinct evolution allows the best resolution in time differences [38]. The unitary transformation performed by the interferometer may, however, allow less distinctness than the energy bound (e.g., for [69] it is half). With monochromatic light and beam splitters, overall transformations are isomorphic to 3D rotations [32, 33, 34] and so all bounds (26) apply. In practice, highest resolution is currently achieved using squeezed states that are not maximally distinct [31, 42]. Their resolution scales sub-linearly with energy, but they can be constructed with macroscopic amounts of energy.
@@
Maximum distinctness defines how distinct states are counted in classical phase space [46, 47], since all width bounds are equivalent to bandwidth bounds for large $N$. Consider the number of distinct shifts allowed by the momentum bandwidth bound (25) for a gas of conserved distinguishable particles: for each spatial degree of freedom, maximum distinctness allows one distinct state per area $h$ of accessible momentum-range $\times$ spatial-length. For identical particles, we divide by a permutation count.
@@
Figure 15: Most-distinct vs Gaussian. Fastest changing states under shifts (Figure 12) give a probability distribution $\mbox{$\left|\psi(x)\right|$}^{2}$ (solid) similar to same-height Gaussian distribution (dashed). Nevertheless, standard-deviation of $x$ for $\mbox{$\left|\psi(x)\right|$}^{2}$ is infinite.
@@
If we model classical mechanics as maximally distinct, though, all evolution is effectively discrete. For large objects, discrete separations $h/2E$ in time can be arbitrarily small, but this is just a property of aggregated change: if enough clocks tick, average time between ticks can be arbitrarily short. Given the energy density $\rho$ in a region, we can identify the scale of individual “ticks” there. If we assume that, at the most microscopic scale, all motion is at the speed of light and all change is motion, we can estimate the distance $\lambda$ between distinct positions as a function of $\rho$. In natural units with $c=1$ and $h=2$, $\rho\,\lambda^{3}=E=p=1/\lambda$, since minimum distinct volume $\lambda^{3}$ must move a width $\lambda$ to move distinctly, so $\lambda=\rho^{-1/4}$. In the extreme case of $\rho$ the Planck density, $\lambda$ is the Planck length and $\tau=1/E=\lambda$ is the Planck time. Effective discreteness provides a fundamental link between classical spacetime and quantum dynamics, and makes discrete integration of Lagrangian action exact [64] (cf. [111, 112, 113, 114]).
@@ Conclusions
Quantum mechanics generalizes Fourier analysis to describe all dynamics as unitary evolution of linear superpositions. Widths of generators of such evolutions play the role of bandwidth in bounding the number of distinct states possible in the course of the evolution. Taking $h$ as the unit of action, any natural width of the eigenvalue distribution, times the length of evolution, normally equals the maximum number of distinct states—up to a factor of order unity. For example, if $\Delta E$ is any deviation-width of the energy distribution for moment one or greater, at most about $\Delta E{\mkern 1.0mu}{\Delta t}$ distinct states can occur in time $\Delta t$}. Action similarly counts distinct states possible in other single-parameter unitary transformations such as a shift in space $(\Delta p_{x}{\mkern 1.0mu}\Delta x)$ or a rotation $(\Delta J_{\phi}\Delta\phi)$, since momentum is the generator of shifts and angular momentum the generator of rotations. Maximum distinctness implies discreteness: continuous evolution becomes just interpolation between distinct states, and discrete integration of expectation values becomes exact. The number of distinct states possible in an evolution defines the resolution with which the length of the evolution can be measured.
@@
Acknowledgments   I thank Gerald J. Sussman and the MIT CSAIL Laboratory for aiding and encouraging this research. I also thank Sussman for valuable discussions, as well as Tom Toffoli, Ted Jacobson, Samuel Braunstein, Lorenzo Maccone, Hal Haggard and Adolfo del Campo.